# Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment

### v3

@@ Abstract
While large-scale unsupervised language models (LMs) capture broad world knowledge and reasoning capabilities, steering their behavior toward desired objectives remains challenging due to the lack of explicit supervision. Existing alignment techniques, such as reinforcement learning from human feedback (RLHF), rely on training a reward model and performing reinforcement learning to align with human preferences. However, RLHF is often computationally intensive, unstable, and sensitive to hyperparameters.
To address these limitations, Direct Preference Optimization (DPO) was introduced as a lightweight and stable alternative, enabling direct alignment of language models with pairwise preference data via classification loss. However, DPO and its extensions generally assume a single static preference distribution, limiting flexibility in multi-objective or dynamic alignment settings.
In this paper, we propose a novel framework: Multi-Preference Lambda-weighted Listwise DPO, which extends DPO to incorporate multiple human preference dimensions (e.g., helpfulness, harmlessness, informativeness) and enables dynamic interpolation through a controllable simplex-weighted formulation. Our method supports both listwise preference feedback and flexible alignment across varying user intents without re-training. Empirical and theoretical analysis demonstrates that our method is as effective as traditional DPO on static objectives while offering greater generality and adaptability for real-world deployment.
@@ Introduction
Large language models (LLMs) trained on large-scale text corpora using next-token prediction objectives have demonstrated strong capabilities across a range of natural language understanding and generation tasks. However, these models often exhibit misalignment with human values and intentions, producing responses that may be factually incorrect, verbose, offensive, or otherwise undesirable. This gap arises in part because the standard unsupervised training objective does not explicitly encode human preferences or task-specific constraints.
To improve alignment with human intent, recent research has introduced fine-tuning techniques that incorporate human feedback. Reinforcement learning from human feedback (RLHF) is one prominent approach: it first trains a reward model from pairwise human comparisons, and then optimizes the language model via reinforcement learning to maximize this reward while maintaining similarity to the original policy. While effective, RLHF is known to be unstable, expensive to train, and sensitive to hyperparameter choices. It also requires online sampling and reward estimation, making it cumbersome to deploy at scale.
Direct Preference Optimization (DPO) was proposed as a simpler alternative to RLHF. Rather than relying on reinforcement learning, DPO directly optimizes the model’s output distribution to reflect human preferences using a pairwise classification objective. DPO has been shown to produce high-quality generations with significantly lower training complexity and resource demands. Despite its advantages, the original DPO formulation is limited in its ability to handle diverse or conflicting user preferences, as it is designed for a single fixed reward distribution.
In practical applications, however, alignment often involves multiple criteria—such as helpfulness, harmlessness, factuality, and succinctness—that may vary in importance depending on the use case or user. This motivates a more flexible framework capable of learning from multiple preference dimensions and generalizing across different preference trade-offs.
In this work, we introduce Multi-Preference Lambda-weighted Listwise DPO, a generalization of the DPO framework that incorporates multiple dimensions of preference feedback and enables interpolation over them via a weight vector $\lambda$ drawn from the probability simplex. Our method extends DPO from pairwise to listwise settings and supports alignment under dynamically changing preference configurations. At inference time, users or downstream systems can modulate the alignment behavior by selecting different $\lambda$ vectors without re-training the model.
We present a detailed theoretical derivation of our method starting from the KL-regularized RL objective, and show how it leads to a clean listwise loss that aggregates multiple reward dimensions. We demonstrate the effectiveness of our approach on multi-objective alignment benchmarks, showing that it achieves competitive or superior results compared to conventional DPO, while offering greater flexibility and control.
@@
Reinforcement Learning from Human Feedback (RLHF).
RLHF [19, 1] has become the de facto paradigm for aligning large language models (LLMs) with human preferences. This framework typically proceeds in three steps: (1) collect human comparisons over model outputs, (2) train a reward model to mimic these preferences, and (3) fine-tune the base model via reinforcement learning, most commonly using Proximal Policy Optimization (PPO) [22]. While effective in improving response helpfulness and safety, RLHF is complex to implement and computationally intensive, requiring online rollouts, heavy sampling, and sensitive hyperparameter tuning [31]. Moreover, it often suffers from reward hacking and policy collapse [4], making it difficult to scale or adapt in practical applications.
Direct Preference Optimization (DPO).
To address the limitations of RLHF, Direct Preference Optimization (DPO) [20] reparameterizes the reward model such that the optimal policy is analytically defined as a function of preference probabilities. This reformulation allows replacing RL with a simple binary classification objective on preference-labeled data. DPO avoids on-policy rollouts, maintains training stability, and significantly reduces computational cost. Several extensions have since been proposed, such as Quark [27] and LLaMA-DPO [16], which demonstrate competitive or superior performance to PPO-based RLHF in tasks like summarization, question answering, and dialogue generation. However, these approaches are limited to pairwise preferences and assume a single fixed alignment objective.
@@
Traditional learning-to-rank methods distinguish between pairwise [3], pointwise [17], and listwise [25] approaches. Listwise methods are known to better capture global ranking structure, and have recently been adopted for preference modeling in LLMs. UltraFeedback [8] introduces a benchmark dataset where human raters assign rankings to multiple outputs per prompt. Listwise approaches in alignment, such as GLOW [26], use these rankings to learn better reward functions. Our work builds upon these insights but optimizes a listwise loss directly in policy space rather than via supervised reward modeling, enabling end-to-end fine-tuning.
Multi-Objective Alignment.
In many alignment scenarios, objectives such as helpfulness, harmlessness, factuality, and informativeness can conflict. Prior work has addressed this challenge via multi-reward reinforcement learning [12], reward decomposition [13], or manually interpolating between multiple reward functions [24]. However, these methods often require training a separate model for each configuration of preference weights. By contrast, our method incorporates preference weight vectors $\lambda$ during training and samples over the probability simplex, enabling the model to generalize to arbitrary trade-offs without retraining. This mechanism draws inspiration from meta-learning and universal value function approximators [21], but is specifically designed for alignment with human judgments.
Controllable and Steerable LMs.
Beyond RLHF, several techniques aim to control model behavior explicitly. These include prompt-based methods such as zero-shot prompting [2], few-shot prompting [30], and chain-of-thought [23], as well as supervised fine-tuning with control tokens [11] or conditioning variables [29]. Constrained decoding [18] and plug-and-play language models [7] provide post-hoc control. However, these methods typically address narrow or static attributes, lacking the flexibility to express complex human preferences. Our method introduces a theoretically grounded and fine-grained approach for multi-dimensional preference steerability without modifying prompts or decoding strategies.
@@
Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one). These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$. A commonly adopted model is the Bradley-Terry (BT) model [bradley1952rank]:
@@
Model comparison: We experiment with two strong open-source language models—LLaMA3-8B and Qwen. These models differ in architecture, pretraining corpora, and tokenizer, providing a robust testbed for evaluating generalizability.
@@
To comprehensively evaluate model behavior, we use six established benchmarks:
  • MMLU [10]: 57 subject areas with around 15K test questions designed to test academic and professional knowledge.
  • ARC (Challenge) [5]: Grade-school level science questions ($\sim$1.1K) focusing on multi-step reasoning.
  • HellaSwag [28]: 10K commonsense sentence completion tasks with subtle distractors.
  • TruthfulQA [15]: 817 questions designed to expose model hallucination and fact inconsistency.
  • Winograd [14]: 273 examples targeting pronoun coreference and syntactic disambiguation.
  • GSM8K [6]: 8.5K arithmetic reasoning problems requiring multi-step solutions.
We evaluate models along several axes:
  • Win Rate (WR): The proportion of pairwise comparisons where the model’s output is preferred by the reward model.
  • lm-eval-harness Score [9]: Average score given by GPT-4 on multi-turn QA prompts, covering helpfulness, correctness, and language quality.
  • Reward Accuracy (%): The percentage of outputs that match the preferred choice of the reward model.
  • KL Divergence: The average KL divergence between the trained policy and the SFT policy, indicating policy shift magnitude.
  • Run Time: Total training duration under fixed hardware settings.
  • Peak GPU Memory: Maximum memory usage during training.
  • Spearman Correlation: Correlation between output log-probability and output length, used to measure length bias.
@@
To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct two complementary sets of ablation experiments, each targeting a different aspect of the framework’s flexibility and effectiveness.
The first set of experiments aims to isolate the effect of explicit preference vector interpolation by comparing $\lambda$-DPO models trained with and without multi-dimensional preference aggregation. In this setting, the random $\lambda$ and Gaussian $\lambda$ variants are used to provide dynamic and smoothly varying interpolation weights over multiple preference dimensions, while the unimodal $\lambda$ variant—represented by fixed one-hot vectors such as $[1,0,0,0]$—serves as the baseline where no interpolation is applied, i.e., each training instance aligns with a single fixed objective. This comparison helps assess whether the simplex-weighted combination of preferences brings measurable benefits in robustness, generalization, or alignment quality.
The second set of experiments directly compares $\lambda$-DPO with the original DPO framework. To ensure a fair comparison that isolates the algorithmic difference—rather than the benefit from multiple objectives—this evaluation uses only the unimodal $\lambda$ setting in $\lambda$-DPO, effectively aligning it with single-objective training, just like standard DPO. This design ensures that both models optimize against one fixed preference dimension at a time, allowing us to accurately assess the impact of listwise supervision and policy reparameterization introduced by $\lambda$-DPO.
For the $\lambda$ schedule, we investigate three variants:
@@
  • Gaussian $\lambda$: we sample each component from a standard normal distribution and then normalize the vector to lie on the simplex, thereby enabling smoothly varying but non-uniform emphasis across dimensions.
  • Unimodal $\lambda$: we use fixed one-hot vectors (e.g., $(1,0,0,\dots)$), effectively disabling interpolation and isolating optimization to a single preference dimension per instance.
Together, these ablations help us understand how balanced weighting, smooth interpolation, and single-objective optimization each contribute to the effectiveness of our listwise $\lambda$-DPO framework.
@@
We conduct a comparative experiment to evaluate the performance of standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups. Specifically, we focus on two representative models—LLaMA 3.2B Instruct and Qwen 1.8B Chat—and assess generalization across five commonly used benchmarks using the lm-eval-harness framework. Both methods are trained from the same base model (i.e., without supervised fine-tuning), using UltraFeedback preference data.
@@
Table 1 reports the accuracy of DPO and $\lambda$-DPO on five representative benchmarks for both LLaMA 3.2 1B Instruct and Qwen 1.8B Chat. On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +1.8% on GSM8K, and +3.6% on TruthfulQA. Qwen 1.8B Chat also sees improvements on most benchmarks, particularly GSM8K (+14.7%) and MMLU (+3.9%). DPO only slightly outperforms $\lambda$-DPO on TruthfulQA in the Qwen setting.
Model Method MMLU ARC HellaSwag TruthfulQA GSM8K
LLaMA 3.2 1B Instruct DPO 0.458 0.364 0.456 0.378 0.696
$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714
Qwen 1.8B Chat DPO 0.379 0.271 0.418 0.366 0.008
$\lambda$-DPO 0.418 0.284 0.428 0.312 0.155
Table 1: Accuracy comparison between DPO and $\lambda$-DPO on five benchmarks using LLaMA 3.2 1B Instruct and Qwen 1.8B Chat. Bold numbers indicate better performance.

### v4

@@ Abstract
While large language models (LLMs) excel at text generation, aligning them with human preferences remains challenging. Reinforcement learning from human feedback (RLHF) improves alignment but is costly and unstable. Direct Preference Optimization (DPO) offers a simpler alternative, yet assumes a fixed, single-dimensional preference.
We propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that supports multiple preference dimensions and dynamic interpolation via a simplex-weighted $\lambda$ vector. Our method enables listwise supervision and flexible alignment without re-training. While our experiments are conducted on 1B–2B scale models, this is an intentional choice: smaller models provide a more stringent testbed where performance improvements more clearly reflect the effectiveness of the alignment strategy itself. Moreover, such models are widely used in compute-constrained applications, making our improvements both methodologically meaningful and practically valuable. Empirical results show that our approach matches or surpasses standard DPO on alignment benchmarks while offering improved adaptability.
@@ Introduction
Large language models (LLMs) have achieved impressive results across natural language tasks, but often misalign with human preferences—producing responses that are factually incorrect, verbose, or unhelpful. This misalignment stems from the next-token prediction objective, which lacks explicit human feedback.
To address this, recent approaches incorporate preference supervision. RLHF trains a reward model from human comparisons and optimizes it via reinforcement learning, but suffers from high cost, instability, and deployment complexity. Direct Preference Optimization (DPO) offers a simpler alternative by directly optimizing model outputs to reflect pairwise preferences, achieving strong results with lower training overhead. However, DPO assumes a single reward dimension and cannot adapt to varying or multi-faceted user intents.
In practice, alignment involves multiple criteria—e.g., helpfulness, honesty, fluency—whose importance may shift across tasks. To address this, we propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that incorporates multiple preference axes and allows dynamic control via a weight vector $\lambda$ drawn from the simplex.
Our method extends DPO from pairwise to listwise supervision and introduces a learnable scheduling mechanism over $\lambda$ vectors, enabling flexible, fine-grained alignment. Empirical results on multi-objective benchmarks show that our method matches or exceeds standard DPO while offering greater control and adaptability.
@@
RLHF and DPO.
Reinforcement learning from human feedback (RLHF) [18, 1] aligns LLMs via reward models and PPO [21], but is costly and unstable [28, 5]. Direct Preference Optimization (DPO) [19] simplifies alignment by reframing it as binary classification on preference pairs, improving stability and efficiency. Extensions like Quark [26] and LLaMA-DPO [16] show strong performance, though limited to fixed, pairwise objectives.
@@
Listwise learning captures global ranking better than pairwise [4, 24]. UltraFeedback [8] provides ranked human feedback, and methods like GLOW [25] train reward models from such data. Our approach differs by optimizing listwise loss directly in policy space, without intermediate reward modeling.
Multi-Objective and Controllable Alignment.
Prior work on multi-reward RL [12], decomposition [13], and reward interpolation [23] requires retraining for each preference trade-off. Our method incorporates a weight vector $\lambda$ during training, enabling generalization to arbitrary preferences. Related ideas include universal value function approximators [20]. Steerable LMs using prompts [3, 22], control tokens [11], or constrained decoding [17] offer limited flexibility. In contrast, our method enables fine-grained, multi-dimensional alignment via $\lambda$ without modifying prompts or decoding.
@@
Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one). These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$. A commonly adopted model is the Bradley-Terry (BT) model [2]:
@@
Model comparison: We experiment with two strong open-source language models—LLaMA and Qwen. These models differ in architecture, pretraining corpora, and tokenizer, providing a robust testbed for evaluating generalizability.
@@
We evaluate model performance using the lm-eval-harness [9] framework, which tests accuracy on a diverse set of benchmarks covering reasoning, factual correctness, and commonsense understanding.
Specifically, we report average accuracy across the following six established datasets:
  • MMLU [10]: 57 subject areas with around 15K test questions targeting academic and professional knowledge.
  • ARC (Challenge) [6]: Grade-school science questions ($\sim$1.1K) requiring multi-step reasoning.
  • HellaSwag [27]: 10K commonsense sentence completion tasks with adversarial distractors.
  • TruthfulQA [15]: 817 questions designed to expose hallucinations and factual errors.
  • Winograd [14]: 273 examples testing pronoun coreference and syntactic disambiguation.
  • GSM8K [7]: 8.5K arithmetic reasoning problems with step-by-step solutions.
@@
To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct ablation experiments focusing on how different preference weightings affect alignment performance.
Specifically, we compare two $\lambda$ schedule variants within the listwise $\lambda$-DPO framework:
@@
  • Unimodal $\lambda$: we use fixed one-hot vectors (e.g., $(1,0,0,\dots)$), effectively disabling interpolation and isolating optimization to a single preference dimension per instance. This variant serves as a control version without multi-dimensional aggregation.
This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting. For a separate comparison between standard DPO and our listwise formulation, please refer to the main benchmark results in Table 1.
@@
We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2B Instruct and Qwen 1.8B Chat. Both models are first supervised fine-tuned (SFT) on UltraChat-200k, then aligned using the UltraFeedback listwise dataset.
Training Details.
We use LoRA (rank 8, all-layer target) with a learning rate of $5\times 10^{-6}$, batch size 1, gradient accumulation steps 8, cosine learning rate schedule, and 5 epochs. The KL penalty $\beta$ is set to 0.1. $\lambda$-DPO uses the lambda_dpo loss with chunk size 1.
Evaluation.
We evaluate on six benchmarks (MMLU, ARC, HellaSwag, TruthfulQA, WSC273, GSM8K) using the lm-eval-harness, reporting accuracy as the primary metric.
@@
Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2B Instruct and Qwen 1.8B Chat. On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +3.3% on GSM8K, and +3.6% on TruthfulQA. The improvement on WSC273 is also evident (+1.8%). For Qwen 1.8B Chat, $\lambda$-DPO outperforms DPO on all benchmarks except TruthfulQA, where DPO achieves a slightly higher score. Notably, large gains are observed on GSM8K (+14.7%) and MMLU (+3.9%), demonstrating the generalization benefits of $\lambda$-DPO across different model families.
Model Method MMLU ARC HellaSwag TruthfulQA WSC273 GSM8K
LLaMA 3.2B Instruct DPO 0.458 0.364 0.456 0.378 0.696 0.312
$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714 0.345
Qwen 1.8B Chat DPO 0.379 0.271 0.418 0.366 0.656 0.008
$\lambda$-DPO 0.418 0.284 0.428 0.312 0.667 0.155
Table 1: Accuracy comparison between DPO and $\lambda$-DPO on six benchmarks using LLaMA 3.2B Instruct and Qwen 1.8B Chat. Bold values indicate better performance for each benchmark.
@@
Ablation Experiment
To further examine the effectiveness of different preference vector settings, we conduct an ablation experiment focusing on four unimodal $\lambda$-DPO models. Each model is trained using a fixed one-hot $\lambda$ vector, targeting only one of the four annotated dimensions: helpfulness, honesty, instruction-following, and truthfulness. This setup allows us to isolate the influence of each individual preference axis and compare their respective impacts across evaluation benchmarks.
Unimodal Comparison Across Benchmarks
We evaluate the four unimodal $\lambda$-DPO variants across six standard benchmarks—mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k. Figure 9 shows the accuracy distribution of each model across these tasks.
Figure 3: Accuracy Comparison of Unimodal $\lambda$-DPO Models with Fixed Preference Vectors.
We observe that all four models achieve relatively comparable performance across the six benchmarks. The $\lambda=[0,1,0,0]$ model (favoring honesty) achieves the highest average accuracy, while the $\lambda=[1,0,0,0]$ model (favoring helpfulness) performs best on truthfulqa and shows strong performance on gsm8k. The models focusing on instruction-following and fluency show only marginally lower scores, with no major degradation in benchmark performance.
Preference Ranking and Vector Hypothesis
To better assess how each individual preference dimension contributes to generalization performance, we compute the average accuracy of the four unimodal $\lambda$-DPO models across six benchmark tasks:
  • $\lambda=[1,0,0,0]$: 0.4593
  • $\lambda=[0,1,0,0]$: 0.4598
  • $\lambda=[0,0,1,0]$: 0.4543
  • $\lambda=[0,0,0,1]$: 0.4548
  • $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$: 0.4612
The results indicate that all four unimodal preference dimensions contribute to model alignment, with helpfulness and honesty yielding slightly better generalization. However, the uniform interpolation $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ achieves the highest overall accuracy (0.4612), outperforming all unimodal variants.
This highlights the benefit of explicitly retaining the $\lambda$ vector for multi-objective supervision. Unlike standard DPO, which implicitly collapses preferences into a fixed form, $\lambda$-DPO provides fine-grained control over alignment trade-offs and leads to more robust generalization.
Discussion and Ablation Analysis
To further understand how individual preference dimensions influence alignment, we observe from Figure 9 that:
  • The $\lambda=[1,0,0,0]$ (helpfulness) model performs best on truthfulqa and achieves solid results on gsm8k.
  • The $\lambda=[0,1,0,0]$ (honesty) model leads slightly in overall average accuracy, especially strong on mmlu.
  • The $\lambda=[0,0,1,0]$ (instruction-following) and $\lambda=[0,0,0,1]$ (fluency) models show close and competitive performance, contrary to the notion that they are stylistically weaker.
These results indicate that optimizing for either content-driven or stylistic objectives in isolation yields comparable outcomes under our listwise framework. It further motivates the integration of multiple preference signals, especially for tasks requiring both correctness and user-aligned expressiveness.
In future work, we plan to extend this ablation framework by exploring:
  • Adaptive $\lambda$ scheduling, where preference weights change based on input context or task type.
  • Hierarchical modeling that separates semantic utility from stylistic fluency.
  • Persona- or task-specific interpolation strategies to match different downstream alignment needs.
Performance-Driven Lambda Scheduling via Polynomial Modeling
We introduce a general and flexible method for performance-driven lambda scheduling, which constructs a probability distribution over preference vectors $\lambda\in\Delta^{d}$ based on observed alignment performance. Each supervised training point $(\lambda^{(i)},y^{(i)})$ consists of a preference configuration $\lambda^{(i)}$ and a corresponding scalar performance metric $y^{(i)}$ (e.g., average benchmark accuracy). The scheduling process unfolds in four steps:
Step 1: Polynomial Feature Mapping
We begin by mapping each preference vector $\lambda$ to a high-dimensional feature space using a polynomial embedding. Specifically, we define a degree-$p$ polynomial feature map $\phi_{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$ that enumerates all monomials up to degree $p$:
$\phi_{p}(\lambda)=\left[1,\lambda_{1},\dots,\lambda_{d},\lambda_{1}^{2},\lambda_{1}\lambda_{2},\dots,\lambda_{d}^{p}\right]^{\top},$ (17)
where the total number of features is given by:
$N=\binom{d+p}{p}.$ (18)
This expansion enables expressive yet tractable modeling of non-linear relationships between $\lambda$ vectors and alignment performance.
Step 2: Polynomial Regression
Using the training set $\{(\lambda^{(i)},y^{(i)})\}_{i=1}^{n}$, we fit a least-squares regression model in the lifted feature space to learn the performance landscape:
$\min_{\mathbf{w}\in\mathbb{R}^{N}}\sum_{i=1}^{n}\left(\mathbf{w}^{\top}\phi_{p}(\lambda^{(i)})-y^{(i)}\right)^{2}.$ (19)
The resulting predictive function
$f(\lambda)=\mathbf{w}^{\top}\phi_{p}(\lambda)$ (20)
approximates the expected performance of any candidate $\lambda$ on downstream alignment tasks.
Step 3: Grid Construction over the Simplex
To enable tractable sampling from the $\lambda$ space, we construct a discrete candidate set $\{\lambda^{(j)}\}_{j=1}^{k}\subset\Delta^{d}$ via one of the following methods:
  • Simplex discretization: generate all combinations with replacement and normalize;
  • Dirichlet sampling: draw $\lambda^{(j)}\sim\mathrm{Dir}(\boldsymbol{\alpha})$, where $\boldsymbol{\alpha}=\mathbf{1}$ yields a uniform prior over the simplex.
Step 4: Softmax-Based Scheduling Distribution
We convert the predicted scores $s_{j}=f(\lambda^{(j)})$ into a probability distribution over the candidate $\lambda$ vectors using a softmax transformation:
$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{k}\exp(\tau\cdot f(\lambda^{(l)}))},$ (21)
where the temperature parameter $\tau>0$ controls the exploration–exploitation trade-off. Higher values of $f(\lambda^{(j)})$ correspond to better expected performance, allowing the scheduler to prioritize promising $\lambda$ vectors while retaining stochasticity. This yields a flexible, differentiable mechanism for guiding multi-objective alignment through learned preference scheduling.
Step 5: Empirical Instantiation with Five Training Points
To validate the polynomial scheduling framework, we use five observed training points—four unimodal $\lambda$ vectors and one uniform configuration—each paired with an empirical alignment accuracy:
$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
We set the polynomial degree to $p=2$, yielding a total of $N=\binom{4+2}{2}=15$ monomial features. Using gradient descent to minimize the squared error over the training set, we obtain the fitted predictive model:
$f(\lambda)$ $=0.35+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
$+0.02\lambda_{1}^{2}+0.04\lambda_{1}\lambda_{2}+0.03\lambda_{1}\lambda_{3}+0.02\lambda_{1}\lambda_{4}$
$+0.02\lambda_{2}^{2}+0.02\lambda_{2}\lambda_{3}+0.02\lambda_{2}\lambda_{4}$
$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
Step 6: Lambda Sampling from Learned Scheduler
We discretize the simplex $\Delta^{4}$ by uniformly sampling $k=10$ candidate $\lambda$ vectors from a Dirichlet$(\boldsymbol{1})$ prior to ensure diverse coverage of the preference space. For each sample $\lambda^{(j)}$, we evaluate its predicted alignment score $s_{j}=f(\lambda^{(j)})$ using the learned polynomial model, and convert these scores into a sampling distribution using softmax:
$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{10}\exp(\tau\cdot f(\lambda^{(l)}))},\quad\text{with }\tau=100.$
This yields a discrete distribution over candidate $\lambda$ vectors that prioritizes those with higher predicted alignment quality. The table below summarizes the 10 sampled vectors, their predicted scores, and their corresponding softmax probabilities:
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $f(\lambda)$ $p(\lambda)$
0.212 0.334 0.245 0.209 0.452 0.095
0.126 0.237 0.131 0.507 0.450 0.075
0.542 0.079 0.256 0.123 0.455 0.117
0.233 0.721 0.020 0.025 0.459 0.185
0.004 0.334 0.281 0.381 0.449 0.067
0.320 0.069 0.513 0.099 0.451 0.082
0.236 0.155 0.358 0.251 0.452 0.093
0.141 0.110 0.701 0.048 0.454 0.110
0.070 0.198 0.388 0.344 0.451 0.081
0.090 0.139 0.306 0.465 0.450 0.095
Table 2: Sampled $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ), their predicted scores $f(\lambda)$, and softmax-based sampling probabilities $p(\lambda)$ with temperature $\tau=100$.
Step 7: Scheduler vs. Uniform Comparison
To assess the overall utility of the learned scheduling distribution, we compare it against a baseline that uses a fixed uniform preference vector $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ for all training examples. Both approaches are applied under identical listwise $\lambda$-DPO training conditions using the LLaMA 3.2B Instruct model.
Figure 4 shows the empirical accuracy across six alignment benchmarks. The learned scheduler achieves a slightly higher average accuracy (0.462 vs. 0.461), validating its effectiveness as a data-driven alternative to static preference weighting.
Beyond average performance, the learned scheduler offers several additional advantages: (i) it introduces sample-level variability by stochastically selecting from a high-quality preference distribution, which can act as an implicit regularizer during training; and (ii) it enables better coverage of the preference simplex, reducing the risk of overfitting to limited preference configurations and helping mitigate modeling bias in the polynomial predictor.
Figure 4: Comparison between learned $\lambda$ scheduler and uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ across six alignment benchmarks. The learned scheduler achieves a slightly higher average accuracy and improves robustness by distributing preference sampling over higher-quality regions of the simplex.